# Sheffer polynomials and the s-ordering of exponential boson operators

### v1

@@ 5 A new parametric family and the main result
The main result (Theorem 5.3) can now be approached. According to Theorem 4.4, the normally ordered ($s=\nobreak-1$) and anti-normally ordered ($s=\nobreak+1$) representations of the exponential operator ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ can be expressed in terms of $\mathcal{HS}(t,z;\allowbreak A,B,r)$, the EGF of the parametric family of exponential Riordan arrays of Hsu–Shiue type, for which there is an explicit expression. (Recall Proposition 3.4.) For any $s\in\mathbb{C}$, the $s$-ordered representation of ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ will now be obtained in terms of a new family of exponential Riordan arrays, or family of Sheffer polynomial sequences, which is of independent interest.
@@ 5.1 The new parametric family
$g_{T}(D;B,r,r^{\prime};\,s)\mathrel{:=}\left(\frac{1+s}{2}\right)\left[\frac{1+(\frac{1-s}{2})B\,D}{1-(\frac{1+s}{2})B\,D}\right]^{-r/B}\!\!+\left(\frac{1-s}{2}\right)\left[\frac{1-(\frac{1+s}{2})B\,D}{1+(\frac{1-s}{2})B\,D}\right]^{r^{\prime}/B},$ (60)
@@
with the cases $A=0$, $B=0$ handled by taking limits. These $g_{T},f_{T}$ have Maclaurin expansions that are series in $\mathbb{C}[[D]]$ of order 0 and $1$ respectively: $g_{T}=1\cdot D^{0}+\nobreak O(D^{1})$ and $f_{T}=1\cdot D^{1}+\nobreak O(D^{2})$.
@@
Here, the pair $\bar{g}_{T},\bar{f}_{T}]$ is the inverse in the exponential Riordan group $\mathfrak{R}$ of the parametric element $g_{T},f_{T}]$ defined in eq. (60). That is, $D=\bar{f}_{T}(z;s)$ is inverse to $z=f_{T}(D;s)$ and can be obtained by series reversion, and $\bar{g}_{T}=1/g_{T}(D;s)=1/g_{T}(\bar{f}_{T}(z;s))$. The notations $\mathbf{S}[g_{T},f_{T}]$ and $\mathbf{R}[\bar{g}_{T},\bar{f}_{T}]$ are equivalent.
It should be clear from the definition of $f_{T}$ in (60) that for general $s$ and $A,B$, the inverse $D=\bar{f}_{T}(z)$ of $z=f_{T}(D)$ is not an elementary function. But:
@@
Set s=1,+1s=-1,+1 in the formulae (60) for $g_{T},f_{T}$, and compare them with the formulae (36) for $d_{\textit{HS}},h_{\textit{HS}}$. The formula for the parametric Hsu–Shiue EGF $\mathcal{HS}(t,z)$ was supplied in Proposition 3.4. ∎
So, exponential Riordan arrays of the newly defined two-point Hsu–Shiue type (and likewise their row polynomials and EGF’s) interpolate as $s$ varies over the $1$-simplex $1\leqslant s\leqslant+1$} between conventional Hsu–Shiue ones, to which they reduce when $s=\pm 1$. One could also define three-point interpolation, with a parameter vector that varies over a $2$-simplex, etc.
@@ 5.2 Main result
The following theorem states how the $s$-ordered representation of ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ can be expressed in terms of the two-point Hsu–Shiue EGF $\mathcal{T}(t,z;A,B,r,r^{\prime};s)$, and subsumes Theorem 4.4, which provided only normal ($s=\nobreak-1$) and anti-normal ($s=\nobreak+1$) orderings. Examples are given in Sec. 6. The corollary to this theorem (Corollary 5.4) expresses $({a^{\dagger}}^{L}a{a^{\dagger}}^{R})^{n}$ in terms of the row polynomials $T_{n}(t;\allowbreak A,B,r,r^{\prime};s)$, and subsumes Theorem 4.2.
@@
${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ $=\mathopen{:}\,\mathrel{\mathcal{T}({a^{\dagger}}a,\lambda{a^{\dagger}}^{e};\,e,1,-L,R;\,s)}\,\mathclose{:}_{s}$ (61)
$=\mathopen{:}\,\mathrel{\bar{g}_{T}(\lambda{a^{\dagger}}^{e};\,s)\exp\left[({a^{\dagger}}a)\bar{f}_{T}(\lambda{a^{\dagger}}^{e};\,s)\right]}\,\mathclose{:}_{s},$ (62)
where $\bar{g}_{T},\bar{f}_{T}]$ is the Riordan-group inverse of $g_{T},f_{T}]$, defined as in eq. (60), with parameters $(A,B,r,r^{\prime};s)=(e,1,-L,R;s)$. This reduces when $s=\pm 1$ to
@@
The idea is this: $G={\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ equals $\mathopen{:}\,\mathrel{F^{(s)}(\alpha,\alpha^{*})}\,\mathclose{:}_{s}$, where $F^{(s)}(\alpha,\alpha^{*})$ is some function that satisfies the diffusion equation (23), a PDE that evolves $F^{(s)}$ toward smaller values of $s$. (In this proof, for clarity the commuting arguments of $F^{(s)}$ will be written as $\alpha,\alpha^{*}$ and not as $a,{a^{\dagger}}$. The notational convention in quantum optics that dependence on $\alpha,\alpha^{*}$ really means dependence on a single complex variable $\alpha=\allowbreak\alpha_{r}+\nobreak{\rm i}\alpha_{i}$ will briefly be followed.)
In principle, one could start with $F^{(+1)}=F^{(A)}$ and integrate the diffusion equation backward in $s$ toward $F^{(-1)}=F^{(N)}$, as one does in quantum optics when smoothing the Glauber–Sudarshan $P$-function ($s=\nobreak+1$) to obtain the Husimi $Q$-function ($s=\nobreak-1$) [2, 3]. The differential operator $\partial^{2}/\partial\alpha\partial\alpha^{*}$ in the PDE is to be interpreted as $\Delta/4$, with $\Delta=\partial^{2}/\partial\alpha_{r}^{2}+\partial^{2}/\partial\alpha_{i}^{2}$ the Laplacian on the complex $\alpha$-plane.
However, to avoid convergence questions the exponential operator $G$ is viewed in this paper as an element of $\mathcal{W}[[\lambda]]$, and $F^{(s)}$ is accordingly an element of $\mathcal{W}_{c}[[\lambda]]$. So $F^{(s)}$ is not a single function on the complex $\alpha$-plane (i.e., a function of $\alpha_{r},\alpha_{i}$ or informally of $\alpha,\alpha^{*}$), and the convolution or smoothing formula (24) cannot easily be used. Instead, $F^{(s)}$ at each value of $s$ is a power series in $\lambda$, with coefficients that are polynomial in $\alpha,\alpha^{*}$, and the series is formal: it is never summed. The following proof should be read with this in mind.
@@
$F^{(s)}(\alpha,\alpha^{*})={V}(\alpha^{*}\alpha,\lambda{\alpha^{*}}^{e};\,s),$ (63)
where $\lambda\neq 0$ and the function $V(t,z;s)$ is to be determined. Changing the independent variables $\alpha,\alpha^{*}$ (really $\alpha_{r},\alpha_{i}$) to $t=\alpha^{*}\alpha$, $z=\lambda{\alpha^{*}}^{e}$ converts the diffusion equation (23) to an equation of diffusion type on the $(t,z)$-plane,
$\left[\frac{\partial}{\partial s}+\mathcal{X}\right]V(t,z;\,s)=0,\qquad\mathcal{X}\mathrel{:=}\frac{1}{2}\left(t\frac{\partial^{2}}{\partial t^{2}}+\frac{\partial}{\partial t}+ez\,\frac{\partial^{2}}{\partial t\,\partial z}\right),$ (64)
@@
$V(\cdot,\cdot;\,s)=\exp\left[(s_{0}-s)\mathcal{X}\right]\,V(\cdot,\cdot;\,s_{0}),\qquad s<s_{0}.$ (65)
@@
$V(t,z;\,s)\varpropto(s_{0}-s)^{-1}{\rm e}^{-2t/(s_{0}-s)},\qquad s<s_{0},$ (66)
for any real $s_{0}$. Depending on $t=\alpha^{*}\alpha$ but not on $z=\lambda{\alpha^{*}}^{e}$, for $s_{0}=+1$ this is in fact the $s$-ordered quasiprobability density on the complex $\alpha$-plane coming from the density operator of the vacuum state [2, §12.4.4], showing how the $P$-function ($s=\nobreak+1$), which is a $\delta$-function located at $\alpha=0$, broadens into a Gaussian as $s$ is decreased and the $Q$-function ($s=\nobreak-1$) is approached. But it is qualitatively different from the expressions $V(t=\alpha^{*}\alpha,\allowbreak z=\lambda{\alpha^{*}}^{e};\pm 1)$ appearing in Theorem 4.4.
@@
$V(t,z;\,s)=\mathbf{R}[\bar{g}(\cdot;s),\bar{f}(\cdot;s)](t,z)=\mathbf{S}[g(\cdot;s),f(\cdot;s)](t,z),$ (67)
@@
$V(t,z;\,s)=\bar{g}(z;\,s)\,{\rm e}^{t\bar{f}(z;\,s)},$ (68)
where as usual, $\bar{g}(z;s)=1\cdot z^{0}+O(z^{1})$ and $D=\bar{f}(z;s)=1\cdot z^{1}+O(z^{2})$, with $g(D;s)=1\cdot D^{0}+O(D^{1})$ and $z=f(D;s)=1\cdot D^{1}+O(D^{2})$. Substituting (68) into (64) and separating the terms proportional to $t^{0}$ and $t^{1}$ gives
$2\bar{g}_{s}+ez\bar{f}\bar{g}_{z}+(\bar{f}+ez\bar{f}_{z})g=0,$ (69a)
$2\bar{f}_{s}+ez\bar{f}\bar{f}_{z}+\bar{f}^{2}=0,$ (69b)
@@
The function (or more accurately formal infinite series) $z=f(D;s)$ is the inverse of $D=\bar{f}(z;s)$, computed by series reversion. With its aid, the independent variables $z,s$ in (69) can be replaced by $D,s$, and the dependent variables $\bar{g},\bar{f}$ by $g,f$, where $g=1/(\bar{g}\circ f)$, i.e., $\bar{g}=1/(g\circ\bar{f})$. This yields the PDEs
$-2f_{D}\,g_{s}+D^{2}f_{D}\,g_{D}+(Df_{D}+ef)g=0,$ (71a)
$-2f_{s}+D^{2}f_{D}+eDf=0,$ (71b)
which are satisfied by $f(D;s)$ and $g(D;s)$. The two-point boundary conditions on these dual functions follow from (70a),(70b) and are
$g(D;\,s=-1)=(1+D)^{-R},$ $g(D;\,s=+1)=(1-D)^{-L},$ (72a)
$f(D;\,s=-1)=\left[1-(1+D)^{-e}\right]/e,$ $f(D;\,s=+1)=\left[(1-D)^{-e}-1\right]/e.$ (72b)
The “dualized” PDEs (71) are linear and of first order, and can be solved by Lagrange’s method of characteristics.
The solution of (71b) is
$z=f(D;\,s)=D^{-e}F(\eta),$ (73)
a similarity solution in which $\eta\mathrel{:=}s-2/D$ is the similarity variable and the function $F(\eta)$ is arbitrary. To match the boundary conditions (72b), choose
$F(\eta)=\frac{\left(\frac{-1-\eta}{2}\right)^{-e}-\left(\frac{1-\eta}{2}\right)^{-e}}{e},$ (74)
@@
$z=f(D;\,s)=\frac{\left[1-(\frac{1+s}{2})D\right]^{-e}-{\left[1+(\frac{1-s}{2})D\right]^{-e}}}{e},$ (75)
with expansion $f(D;s)=1\cdot D^{1}+O(D^{2})$. With this choice of f(D;s)f(D;\,s), the PDE (71a) is solved in the same way. It has the similarity solution
$g(D;\,s)=\left[\left(\frac{1+s}{2}\right)+\left(\frac{1-s}{2}\right)\left(\frac{\eta+1}{\eta-1}\right)^{e+1}\right]G(\eta),$ (76)
where the function $G(\eta)$ is arbitrary. To match the boundary conditions (72a), choose
$G(\eta)=\left(\frac{\eta-1}{\eta+1}\right)^{L},$ (77)
@@
$g(D;\,s)=\left(\frac{1+s}{2}\right)\left[\frac{1+(\frac{1-s}{2})D}{1-(\frac{1+s}{2})D}\right]^{L}+\left(\frac{1-s}{2}\right)\left[\frac{1-(\frac{1+s}{2})D}{1+(\frac{1-s}{2})D}\right]^{R},$ (78)
with $g(D;s)=1\cdot D^{0}+O(D^{1})$. Comparing (75),(78) with the definitions of $f_{T},g_{T}$ given in (60) reveals that
$f(D;\,s)=f_{T}(D;e,1;\,s),\qquad g(D;\,s)=g_{T}(D;1,-L,R;\,s).$ (79)
@@
Wilcox [49] solves a PDE similar to (64) by normal ordering an exponentiated second-order differential operator analogous to the $\exp\left[(s_{0}-\nobreak s)\mathcal{X}\right]$ in eq. (65). In his paper the PDE comes first, but in the just-concluded proof of Theorem 5.3 the normal ordering (actually $s$-ordering) came first, and determined the PDE to be solved.
@@
$({a^{\dagger}}^{L}a{a^{\dagger}}^{R})^{n}=\mathopen{:}\,\mathrel{T_{n}({a^{\dagger}}a;e,1,-L,R;s)}\,\mathclose{:}_{s},$
where $T_{n}(t;e,1,-L,R;s)$ is the case $(A,B,r,r^{\prime})=(e,1,-L,R)$ of the nn\!’th parametric Sheffer polynomial $\mathbf{S}[g_{T},f_{T}]_{n}(t)=\mathbf{R}[\bar{g}_{T},\bar{f}_{T}]_{n}(t)$, with $g_{T}(D;s)$ and $z=f_{T}(D;s)$ given in Definition 5.1, i.e., in eq. (60).
@@
Expand each side of the identity (61) in Theorem 5.3 as a formal power series in $\lambda$, and equate the coefficients of $\lambda^{n}$ on the two sides. ∎
The following is the Weyl ordering ($s=\nobreak 0$) case of the main result Theorem 5.3, which is the one that can be expressed most simply. The functions $\bar{g}(z)$, $D=\bar{f}(z)$ come from $g(D)$, $z=f(D)$, which in turn come from the two-point functions $g_{T}(D;s)$, $z=f_{T}(D;s)$ defined in (60), by setting $s=\nobreak 0$.

### v2

@@ 5 A new parametric family and the main result
The main result (Theorem 5.3) can now be approached. According to Theorem 4.4, the normally ordered ($s=\penalty 10000-1$) and anti-normally ordered ($s=\penalty 10000+1$) representations of the exponential operator ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ can be expressed in terms of $\mathcal{HS}(t,z;\allowbreak A,B,r)$, the EGF of the parametric family of exponential Riordan arrays of Hsu–Shiue type, for which there is an explicit expression. (Recall Proposition 3.4.) For any $s\in\mathbb{C}$, the $s$-ordered representation of ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ will now be obtained in terms of a new family of exponential Riordan arrays, or family of Sheffer polynomial sequences, which is of independent interest.
@@ 5.1 The new parametric family
$g_{T}(D;B,r,r^{\prime};\,s)\mathrel{:=}\left(\frac{1+s}{2}\right)\left[\frac{1+(\frac{1-s}{2})B\,D}{1-(\frac{1+s}{2})B\,D}\right]^{-r/B}\!\!+\left(\frac{1-s}{2}\right)\left[\frac{1-(\frac{1+s}{2})B\,D}{1+(\frac{1-s}{2})B\,D}\right]^{r^{\prime}/B},$ (5.1)
@@
with the cases $A=0$, $B=0$ handled by taking limits. These $g_{T},f_{T}$ have Maclaurin expansions that are series in $\mathbb{C}[[D]]$ of order 0 and $1$ respectively: $g_{T}=1\cdot D^{0}+\penalty 10000O(D^{1})$ and $f_{T}=1\cdot D^{1}+\penalty 10000O(D^{2})$.
@@
Here, the pair $\bar{g}_{T},\bar{f}_{T}]$ is the inverse in the exponential Riordan group $\mathfrak{R}$ of the parametric element $g_{T},f_{T}]$ defined in eq. (5.1). That is, $D=\bar{f}_{T}(z;s)$ is inverse to $z=f_{T}(D;s)$ and can be obtained by series reversion, and $\bar{g}_{T}=1/g_{T}(D;s)=1/g_{T}(\bar{f}_{T}(z;s))$. The notations $\mathbf{S}[g_{T},f_{T}]$ and $\mathbf{R}[\bar{g}_{T},\bar{f}_{T}]$ are equivalent.
It should be clear from the definition of $f_{T}$ in (5.1) that for general $s$ and $A,B$, the inverse $D=\bar{f}_{T}(z)$ of $z=f_{T}(D)$ is not an elementary function. But:
@@
Set s=1,+1s=-1,+1 in the formulae (5.1) for $g_{T},f_{T}$, and compare them with the formulae (3.13) for $d_{\textit{HS}},h_{\textit{HS}}$. The formula for the parametric Hsu–Shiue EGF $\mathcal{HS}(t,z)$ was supplied in Proposition 3.4. ∎
So, exponential Riordan arrays of the newly defined two-point Hsu–Shiue type (and likewise their row polynomials and EGF’s) interpolate as the parameter $s$ varies over the $1$-simplex $1\leqslant s\leqslant+1$} between a pair of conventional Hsu–Shiue ones, to which they reduce when $s=\pm 1$. One could also define three-point interpolation, with a parameter vector that varies over a $2$-simplex, etc.
@@ 5.2 Main result
The following theorem states how the $s$-ordered representation of ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ can be expressed in terms of the two-point Hsu–Shiue EGF $\mathcal{T}(t,z;A,B,r,r^{\prime};s)$, and subsumes Theorem 4.4, which provided only normal ($s=\penalty 10000-1$) and anti-normal ($s=\penalty 10000+1$) orderings. Examples are given in Sec. 6. The corollary to this theorem (Corollary 5.4) expresses $({a^{\dagger}}^{L}a{a^{\dagger}}^{R})^{n}$ in terms of the row polynomials $T_{n}(t;\allowbreak A,B,r,r^{\prime};s)$, and subsumes Theorem 4.2.
@@
${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ $=\mathopen{:}\,\mathrel{\mathcal{T}({a^{\dagger}}a,\lambda{a^{\dagger}}^{e};\,e,1,-L,R;\,s)}\,\mathclose{:}_{s}$ (5.2)
$=\mathopen{:}\,\mathrel{\bar{g}_{T}(\lambda{a^{\dagger}}^{e};\,s)\exp\left[({a^{\dagger}}a)\bar{f}_{T}(\lambda{a^{\dagger}}^{e};\,s)\right]}\,\mathclose{:}_{s},$ (5.3)
where $\bar{g}_{T},\bar{f}_{T}]=[\bar{g}_{T}(z;s),\bar{f}_{T}(z;s)]$ is the Riordan-group inverse of $g_{T},f_{T}]$, defined as in eq. (5.1), with parameters $(A,B,r,r^{\prime};s)=(e,1,-L,R;s)$. This reduces when $s=\pm 1$ to
@@
The idea is this: $G={\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ equals $\mathopen{:}\,\mathrel{F^{(s)}(\alpha,\alpha^{*})}\,\mathclose{:}_{s}$, where $F^{(s)}(\alpha,\alpha^{*})$ is some function that satisfies the diffusion equation (2.19), a PDE that evolves $F^{(s)}$ toward smaller values of $s$. (In this proof, for clarity the commuting arguments of $F^{(s)}$ will be written as $\alpha,\alpha^{*}$ and not as $a,{a^{\dagger}}$. The notational convention in quantum optics that dependence on $\alpha,\alpha^{*}$ really means dependence on a single complex variable $\alpha=\allowbreak\alpha_{r}+\penalty 10000{\rm i}\alpha_{i}$ will briefly be followed.)
In principle, one could start with $F^{(+1)}=F^{(A)}$ and integrate the diffusion equation backward in $s$ toward $F^{(-1)}=F^{(N)}$, as one does in quantum optics when smoothing the Glauber–Sudarshan $P$-function ($s=\penalty 10000+1$) to obtain the Husimi $Q$-function ($s=\penalty 10000-1$) [2, 3]. The differential operator $\partial^{2}/\partial\alpha\partial\alpha^{*}$ in the PDE is to be interpreted as $\Delta/4$, with $\Delta=\partial^{2}/\partial\alpha_{r}^{2}+\partial^{2}/\partial\alpha_{i}^{2}$ the Laplacian on the complex $\alpha$-plane.
However, to avoid convergence questions the exponential operator $G$ is viewed in this paper as an element of $\mathcal{W}[[\lambda]]$, and $F^{(s)}$ is accordingly an element of $\mathcal{W}_{c}[[\lambda]]$. So $F^{(s)}$ is not a single function on the complex $\alpha$-plane (i.e., a function of $\alpha_{r},\alpha_{i}$ or informally of $\alpha,\alpha^{*}$), and the convolution or smoothing formula (2.20) cannot easily be used. Instead, $F^{(s)}$ at each value of $s$ is a power series in $\lambda$, with coefficients that are polynomial in $\alpha,\alpha^{*}$, and the series is formal: it is never summed. The following proof should be read with this in mind.
@@
$F^{(s)}(\alpha,\alpha^{*})={V}(\alpha^{*}\alpha,\lambda{\alpha^{*}}^{e};\,s),$ (5.4)
where $\lambda\neq 0$ and the function $V(t,z;s)$ is to be determined. Changing the independent variables $\alpha,\alpha^{*}$ (really $\alpha_{r},\alpha_{i}$) to $t=\alpha^{*}\alpha$, $z=\lambda{\alpha^{*}}^{e}$ converts the diffusion equation (2.19) to an equation of diffusion type on the $(t,z)$-plane,
$\left[\frac{\partial}{\partial s}+\mathcal{X}\right]V(t,z;\,s)=0,\qquad\mathcal{X}\mathrel{:=}\frac{1}{2}\left(t\frac{\partial^{2}}{\partial t^{2}}+\frac{\partial}{\partial t}+ez\,\frac{\partial^{2}}{\partial t\,\partial z}\right),$ (5.5)
@@
$V(\cdot,\cdot;\,s)=\exp\left[(s_{0}-s)\mathcal{X}\right]\,V(\cdot,\cdot;\,s_{0}),\qquad s<s_{0}.$ (5.6)
@@
$V(t,z;\,s)\varpropto(s_{0}-s)^{-1}{\rm e}^{-2t/(s_{0}-s)},\qquad s<s_{0},$ (5.7)
for any real $s_{0}$. Depending on $t=\alpha^{*}\alpha$ but not on $z=\lambda{\alpha^{*}}^{e}$, for $s_{0}=+1$ this is in fact the $s$-ordered quasiprobability density on the complex $\alpha$-plane coming from the density operator of the vacuum state [2, §12.4.4], showing how the $P$-function ($s=\penalty 10000+1$), which is a $\delta$-function located at $\alpha=0$, broadens into a Gaussian as $s$ is decreased and the $Q$-function ($s=\penalty 10000-1$) is approached. But it is qualitatively different from the expressions $V(t=\alpha^{*}\alpha,\allowbreak z=\lambda{\alpha^{*}}^{e};\allowbreak s=\pm 1)$ appearing in Theorem 4.4.
@@
$V(t,z;\,s)=\mathbf{R}[\bar{g}(\cdot;s),\bar{f}(\cdot;s)](t,z)=\mathbf{S}[g(\cdot;s),f(\cdot;s)](t,z),$ (5.8)
@@
$V(t,z;\,s)=\bar{g}(z;\,s)\,{\rm e}^{t\bar{f}(z;\,s)},$ (5.9)
where as usual, $\bar{g}(z;s)=1\cdot z^{0}+O(z^{1})$ and $D=\bar{f}(z;s)=1\cdot z^{1}+O(z^{2})$, with $g(D;s)=1\cdot D^{0}+O(D^{1})$ and $z=f(D;s)=1\cdot D^{1}+O(D^{2})$. Substituting (5.9) into (5.5) and separating the terms proportional to $t^{0}$ and $t^{1}$ gives
$2\bar{g}_{s}+ez\bar{f}\bar{g}_{z}+(\bar{f}+ez\bar{f}_{z})g=0,$ (5.10a)
$2\bar{f}_{s}+ez\bar{f}\bar{f}_{z}+\bar{f}^{2}=0,$ (5.10b)
@@
The function (or more accurately formal infinite series) $z=f(D;s)$ is the inverse of $D=\bar{f}(z;s)$, computed by series reversion. With its aid, the independent variables $z,s$ in (5.10) can be replaced by $D,s$, and the dependent variables $\bar{g},\bar{f}$ by $g,f$, where $g=1/(\bar{g}\circ f)$, i.e., $\bar{g}=1/(g\circ\bar{f})$. This yields the PDEs
$-2f_{D}\,g_{s}+D^{2}f_{D}\,g_{D}+(Df_{D}+ef)g=0,$ (5.12a)
$-2f_{s}+D^{2}f_{D}+eDf=0,$ (5.12b)
which are satisfied by $f(D;s)$ and $g(D;s)$. The two-point boundary conditions on these dual functions follow from (5.11a),(5.11b) and are
$g(D;\,s=-1)=(1+D)^{-R},$ $g(D;\,s=+1)=(1-D)^{-L},$ (5.13a)
$f(D;\,s=-1)=\left[1-(1+D)^{-e}\right]/e,$ $f(D;\,s=+1)=\left[(1-D)^{-e}-1\right]/e.$ (5.13b)
The “dualized” PDEs (5.12) are linear and of first order, and can be solved by Lagrange’s method of characteristics.
The solution of (5.12b) is
$z=f(D;\,s)=D^{-e}F(\eta),$ (5.14)
a similarity solution in which $\eta\mathrel{:=}s-2/D$ is the similarity variable and the function $F(\eta)$ is arbitrary. To match the boundary conditions (5.13b), choose
$F(\eta)=\frac{\left(\frac{-1-\eta}{2}\right)^{-e}-\left(\frac{1-\eta}{2}\right)^{-e}}{e},$ (5.15)
@@
$z=f(D;\,s)=\frac{\left[1-(\frac{1+s}{2})D\right]^{-e}-{\left[1+(\frac{1-s}{2})D\right]^{-e}}}{e},$ (5.16)
with expansion $f(D;s)=1\cdot D^{1}+O(D^{2})$. With this choice of f(D;s)f(D;\,s), the PDE (5.12a) is solved in the same way. It has the similarity solution
$g(D;\,s)=\left[\left(\frac{1+s}{2}\right)+\left(\frac{1-s}{2}\right)\left(\frac{\eta+1}{\eta-1}\right)^{e+1}\right]G(\eta),$ (5.17)
where the function $G(\eta)$ is arbitrary. To match the boundary conditions (5.13a), choose
$G(\eta)=\left(\frac{\eta-1}{\eta+1}\right)^{L},$ (5.18)
@@
$g(D;\,s)=\left(\frac{1+s}{2}\right)\left[\frac{1+(\frac{1-s}{2})D}{1-(\frac{1+s}{2})D}\right]^{L}+\left(\frac{1-s}{2}\right)\left[\frac{1-(\frac{1+s}{2})D}{1+(\frac{1-s}{2})D}\right]^{R},$ (5.19)
with $g(D;s)=1\cdot D^{0}+O(D^{1})$. Comparing (5.16),(5.19) with the definitions of $f_{T},g_{T}$ given in (5.1) reveals that
$f(D;\,s)=f_{T}(D;e,1;\,s),\qquad g(D;\,s)=g_{T}(D;1,-L,R;\,s).$ (5.20)
@@
Wilcox [50] solved a PDE similar to (5.5) by normal ordering an exponentiated second-order differential operator analogous to the $\exp\left[(s_{0}-\penalty 10000s)\mathcal{X}\right]$ in eq. (5.6). In his paper the PDE came first, but in the just-concluded proof of Theorem 5.3 the normal ordering (actually $s$-ordering) came first, and determined the PDE to be solved.
@@
$({a^{\dagger}}^{L}a{a^{\dagger}}^{R})^{n}=\mathopen{:}\,\mathrel{\left({a^{\dagger}}^{e}\right)^{n}T_{n}({a^{\dagger}}a;e,1,-L,R;s)}\,\mathclose{:}_{s},$
where $T_{n}(t;e,1,-L,R;s)$ is the case $(A,B,r,r^{\prime})=(e,1,-L,R)$ of the nn\!’th parametric Sheffer polynomial $\mathbf{S}[g_{T},f_{T}]_{n}(t)=\mathbf{R}[\bar{g}_{T},\bar{f}_{T}]_{n}(t)$, with $g_{T}(D;s)$ and $z=f_{T}(D;s)$ given in Definition 5.1, i.e., in eq. (5.1).
@@
Expand each side of the identity (5.2) in Theorem 5.3 as a formal power series in $\lambda$, and equate the coefficients of $\lambda^{n}$ on the two sides. ∎
The s=1s=-1 (normal ordering) case of the formula in this corollary is an illustration of Prop. 4.1, the result of Duchamp et al.
[…]